OpenStudy (malaya):

V=πr2h Using the formula find h when V = 141.3 and r = 3. Use pi = 3.14.

OpenStudy (fortytherapper):

Have you begun plugging in points?

OpenStudy (malaya):

nope

OpenStudy (fortytherapper):

How would it look if you plugged in all given information V = (pi)(r)^2(h) I'll start it off Since V = 141.3 141.3 = (pi)(r)^2(h) Can you finish it off?

OpenStudy (wolf1728):

141.3 = 3.14159 * 3 * 3 * height CAn you solve that for height?

OpenStudy (malaya):

I don't even know how to do this tbh : / @wolf1728

OpenStudy (wolf1728):

Well, at least you could do the multiplication. Okay I'll do it for you. 141.3 = 28.27433 * height can you solve for height?

OpenStudy (wolf1728):

height = 141.3 / 28.27433

OpenStudy (malaya):

Soo it would be 5.05 ? @wolf1728

OpenStudy (wolf1728):

No - divide 141.3 by 28.27433

OpenStudy (malaya):

Well I put 5.05 and it said it was right but thats what i did @wolf1728

OpenStudy (wolf1728):

Just to make sure, in the problem, does it state the volume is 141.3 and radius is 3?

OpenStudy (malaya):

yeahh

OpenStudy (wolf1728):

Hmmm if the volume actually is 143.1 then the answer comes out to 5.06

OpenStudy (malaya):

its 141.3 @wolf1728

OpenStudy (wolf1728):

Okay, then I don't know how the answer comes to 5.05. You are determining the volume of a cylinder right?

OpenStudy (malaya):

Nope it just says find h @wolf1728

OpenStudy (wolf1728):

Well, whatever the case, you found the correct answer!

OpenStudy (malaya):

Lol yeahh Thankss : )) @wolf1728

OpenStudy (wolf1728):

Okay - well glad to help!